OpenStudy (anonymous):

bobby had b marbles and charlie had c marbles.After bobby gave 6 marbles to charlie,bobby still had 18 more marbles than charlie. find c-b a) 30 b) 12 c) 3 d) -12 e) -30

OpenStudy (ddcamp):

Bobby gives away six. At the end, he has b-6 marbles. Charlie gains six. At the end, he has c+6 marbles. b-6 is 18 greater than c+6. Written as an equation: b-6 = c+6 + 18 b-6 = c + 24 You can solve for (c-b) by subtracting 24 from both sides and subtracting b from both sides: c-b = -24-6 c-b = -30; e

OpenStudy (anonymous):

@DDCamp thanx